Suppose a small finite population consists of only N = 5 numbers:

54 55 59 63 64

I. Find population mean and standard deviation
II. Draw all possible samples of size 2 with replacement
III. Construct sampling distribution
IV. Find sample mean and standard deviation
V. Verify/justify your results.
1
Expert's answer
2021-10-13T11:54:29-0400

I.


"mean=\mu=\dfrac{54 +55 +59+ 63 +64}{5}=59"

"variance=\sigma^2=\dfrac{1}{5}((54-59)^2+(55-59)^2"

"+(59-59)^2+(63-59)^2+(64-59)^2)=16.4"

"\sigma=\sqrt{\sigma^2}=\sqrt{16.4}=2\sqrt{4.1}"

II. There are "5^2=25" samples of size two which can be drawn with replacement: 


"\begin{matrix}
 Sample & Sample\ mean \\
 (54,54) & 54 \\
 (54,55) & 54.5 \\
 (54,59) & 56.5 \\
 (54,63) & 58.5 \\
 (54,64) & 59 \\
 (55,54) & 54.5 \\
 (55,55) & 55\\
 (55,59) & 57 \\
 (55,63) & 59 \\
 (55,64) & 59.5 \\
 (59,54) & 56.5 \\
 (59,55) & 57 \\
 (59,59) & 59 \\
 (59,63) & 61 \\
 (59,64) & 61.5 \\
 (63,54) & 58.5 \\
 (63,55) & 59\\
 (63,59) & 61 \\
 (63,63) & 63 \\
 (63,64) & 63.5 \\
 (64,54) & 59 \\
 (64,55) & 59.5 \\
 (64,59) & 61.5 \\
 (64,63) & 63.5 \\
 (64,64) & 64 \\
\end{matrix}"

III.



"\begin{matrix}
 \bar{X} & P(\bar{X}) \\
 54 & 1/25 \\
 54.5 & 2/25 \\
 55 & 1/25 \\
 56.5 & 2/25 \\
 57 & 2/25 \\
 58.5 & 2/25 \\
 59 & 5/25\\
 59.5 & 2/25 \\
 61 & 2/25 \\
 61.5 & 2/25 \\
 63 & 1/25 \\
 63.5 & 2/25 \\
 64 & 1/25 \\
\end{matrix}"

IV.


"\mu_{\bar{X}}=54(\dfrac{1}{25})+54.5(\dfrac{2}{25})+55(\dfrac{1}{25})+56.5(\dfrac{2}{25})"

"+57(\dfrac{2}{25})+58.5(\dfrac{2}{25})+59(\dfrac{5}{25})+59.5(\dfrac{2}{25})"

"+61(\dfrac{2}{25})+61.5(\dfrac{2}{25})+63(\dfrac{1}{25})+63.5(\dfrac{2}{25})"

"+64(\dfrac{1}{25})=59"


"\sum_i\bar{X}_i^2P(\bar{X_i})=54^2(\dfrac{1}{25})+54.5^2(\dfrac{2}{25})+55^2(\dfrac{1}{25})"

"+56.5^2(\dfrac{2}{25})+57^2(\dfrac{2}{25})+58.5^2(\dfrac{2}{25})+59^2(\dfrac{5}{25})"

"+59.5^2(\dfrac{2}{25})+61^2(\dfrac{2}{25})+61.5^2(\dfrac{2}{25})+63^2(\dfrac{1}{25})"

"+63.5^2(\dfrac{2}{25})+64^2(\dfrac{2}{25})=3489.2"

"\sigma_{\bar{X}}^2=\sum_i\bar{X}_i^2P(\bar{X_i})-\mu_{\bar{X}}^2=3489.2-59^2=8.2"

"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{8.2}"

"\mu_{\bar{X}}=59, \sigma_{\bar{X}}=\sqrt{8.2}"

V.

The mean "\mu_{\bar{X}}" and standard deviation "\sigma_{\bar{X}}" of the sample mean "\bar{X}" satisfy


"\mu_{\bar{X}}=59=\mu, \sigma_{\bar{X}}=\sqrt{8.2}=\dfrac{2\sqrt{4.1}}{\sqrt{2}}=\dfrac{\sigma}{\sqrt{n}}"